Battery A 12.6 volts with an internal resistance of 0.2 ohm and battery B 12.2 volts with an internal resistance of 0.3 ohm are delivering a current through the load RL=10 ohms. Find the voltage across the load.


1
Expert's answer
2022-01-31T10:24:34-0500

∆V=12.6-12.4=0.2V

Load current


"I_L=\frac{∆V}{R_1+R_2}=\frac{0.2}{0.2+0.3}=0.04A"

"Voltage across load \\
∆V=I_LR_L\\∆V=0.04\times10=0.4V"